OpenStudy (anonymous):

Anyone who took Pre-Calculus and understands it, please help me. I'm drowning.

OpenStudy (e.mccormick):

Did you have a question that you needed help with?

OpenStudy (anonymous):

Try two whole topics that I need help with. But if it needs to be narrowed down to a question, how about this. Solve the equation on the interval [0,2pi) \[\cos2x=\sqrt{3}/2\]

OpenStudy (e.mccormick):

So what part of that are you having trouble with?

OpenStudy (anonymous):

Well, you see, our teacher hasn't covered this or reviewed it since...2nd quarter? And now it's on our final and I don't understand how to do any of the identities or logarithms. So...basically I'm having trouble with the whole equation.

OpenStudy (e.mccormick):

Well, what about the unit circle?

OpenStudy (anonymous):

I don't remember how to use the unit circle. Do I use it for this? And if so, how?

OpenStudy (e.mccormick):

Lets start with let \(\theta=2x\) for a quick substitution. That will have to go backwards later, but it is for a reason. Now, look at the unit circle: http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg Where does \(\cos\theta = \frac{\sqrt{3}}{2}\) ?

OpenStudy (anonymous):

\[\pi/6\] and \[11\pi/6\]

OpenStudy (e.mccormick):

Yes. Now, that was for \(\theta\), not for x. So, we have to convert it into x. Also, your x can be between \([0,2\pi)\), so there may be more answers!

OpenStudy (e.mccormick):

So, if we used \(\theta=2x\), what does that make those first two points in terms of x?

OpenStudy (anonymous):

Doubled?

OpenStudy (e.mccormick):

\(\theta=2x\implies \frac{1}{2}\cdot\theta=\frac{1}{2}\cdot2x\)....

OpenStudy (anonymous):

Oh, so it would be half?

OpenStudy (e.mccormick):

Yes. So, what is half of those first two points? Those we know will be answers.

OpenStudy (anonymous):

Ok. That would be... \[22\pi/6\] and \[2\pi/6\]

OpenStudy (anonymous):

No...that's wrong.

OpenStudy (e.mccormick):

You did the 2, not the half. Hehe. Yah. 6 becomes 12...

OpenStudy (anonymous):

\[\pi/12\] and \[11\pi/12\]

OpenStudy (e.mccormick):

So put \(\frac{\sqrt{\pi}}{12}\) and \(\frac{\sqrt{11\pi}}{12}\) aside as answers. We know those are between \([0,2\pi)\). The next thing is to see if there are more. Now, the basis we found was the points \(\frac{\sqrt{\pi}}{6}\) and \(\frac{\sqrt{11\pi}}{6}\). If there were more points, they would be co-terminal angles of those. What would be the next co-terminal angle for each of those? And, is half of the next co-terminal angle \(\ge 2\pi\)?

OpenStudy (e.mccormick):

Aak... I don't know why I put the root in there.. . LOL

OpenStudy (anonymous):

Haha

OpenStudy (e.mccormick):

OH, copied the text and pasted... and forgot to remove the root. LOL

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Could you remind me what a co-terminal angle is? Isn't it an angle on the opposite side of the unit circle?

OpenStudy (e.mccormick):

So put \(\frac{\pi}{12}\) and \(\frac{11\pi}{12}\) aside as answers. We know those are between \([0,2\pi)\). The next thing is to see if there are more. Now, the basis we found was the points \(\frac{\pi}{6}\) and \(\frac{11\pi}{6}\). If there were more points, they would be coterminal angles of those. What would be the next coterminal angle for each of those? And, is half of the next coterminal angle \(\ge 2\pi\)? That is more like it! Ignore the man behind the curtain!

OpenStudy (anonymous):

Hahaha! Nice reference.

OpenStudy (e.mccormick):

co means same, terminal means end.

OpenStudy (anonymous):

I know what the word means. I'm saying what does it mean in the context of Pre-Calc?

OpenStudy (e.mccormick):

Exactly that. Angles that have the same end point.

OpenStudy (anonymous):

Oh. Okay. Haha.

OpenStudy (e.mccormick):

So \(\pm360^\circ\) or \(\pm 2\pi\), with next being the +.

OpenStudy (anonymous):

Oh! I remember that! Ok.

OpenStudy (anonymous):

Would it be \[23\pi/6\] and \[13\pi/6\]

OpenStudy (e.mccormick):

\[\frac{\pi}{6}+\frac{12\pi}{6}=\frac{13\pi}{6}\implies \frac{13\pi}{12}\] \[\frac{11\pi}{6}+\frac{12\pi}{6}=\frac{23\pi}{6}\implies \frac{23\pi}{12}\]

OpenStudy (anonymous):

Wait, how did you get the denominator to 12?

OpenStudy (e.mccormick):

I am just showing the x version. Remember, we still do the theta to x on these. Now, the stopping point is \([0,2\pi)\). Either of those over that line?

OpenStudy (anonymous):

Ummm...Yes. Both of them are.

OpenStudy (e.mccormick):

What is \(2\pi\) when done as a fraction over 12?

OpenStudy (anonymous):

Ohhh...I'm tired, not thinking straight. Lol sorry. \[24\pi/12\]

OpenStudy (e.mccormick):

All I did in one step was this stuff:\[\theta =\frac{13\pi}{6}\implies 2x =\frac{13\pi}{6}\implies \frac{1}{2}\cdot 2x =\frac{1}{2}\cdot \frac{13\pi}{6}\implies x =\frac{13\pi}{12}\] Didn't mean to catch you off guard. Correct, \(2\pi=\frac{24\pi}{12}\). So, are either of those two answers \(\ge \frac{24\pi}{12}\)

OpenStudy (anonymous):

No.

OpenStudy (e.mccormick):

So those are withn the domain and therefore answers.

OpenStudy (e.mccormick):

Now we have 4 answers. So, based on the highest answer, do you think we need to search for more?

OpenStudy (anonymous):

No. We have enough. This test is multiple choice, by the way. And one of the answers is \[\pi/12, 11\pi/12, 13\pi/12, 23\pi/12\]

OpenStudy (anonymous):

So, we're good! :-)

OpenStudy (e.mccormick):

In formal notation, the answer is: \[x=\left \{\frac{\pi}{12},\frac{11\pi}{12},\frac{13\pi}{12},\frac{23\pi}{12}\right \}\]which you don't need for multiple choice...

OpenStudy (anonymous):

Yeah, lol Ok. Now the next one looks confusing too, if it's not too much trouble.

OpenStudy (e.mccormick):

If you look at the sequence, the next one would have been \(\frac{25\pi}{12}\). Comes in handy if you need the answers as a middle part of something else. Find the sequence, and then they are easy to see. So what is the next one?

OpenStudy (anonymous):

Thank you so much, by the way. :)

OpenStudy (anonymous):

The next one is...

OpenStudy (anonymous):

\[\tan 2x-\tan x=0\]

OpenStudy (e.mccormick):

OK, so you need to get that to one angle. Right now you have two angles, 2x and x.

OpenStudy (anonymous):

So how would I go about that?

OpenStudy (e.mccormick):

Have a list of trig identities?

OpenStudy (anonymous):

Ummm...yes I do. Lemme grab it.

OpenStudy (anonymous):

Got it.

OpenStudy (e.mccormick):

We want something that changes \(\tan 2x\) into \(\tan x\) or the other way around.

OpenStudy (anonymous):

Well, there's the Double-Angle Formulas.

OpenStudy (e.mccormick):

Yep! And tan is the ugly one, but oh well.

OpenStudy (anonymous):

\[\tan 2u=2\tan u/1-\tan ^{2}u\]

OpenStudy (e.mccormick):

Ok, so there goes 2x. Assassinated and replaced with something "friendlier" to manipulation.

OpenStudy (anonymous):

Lol good.

OpenStudy (e.mccormick):

So, you have a fraction minus a something. Common denominator time.

OpenStudy (anonymous):

I feel so dumb. What do I to get the common denominator? Can I just put the tanx on the denominator of the other one?

OpenStudy (e.mccormick):

You must use 1 in disguise. Remember that for any value, equation, etc. \(\frac{n}{n}=1\)

OpenStudy (anonymous):

So, I put the tanx over a 1?

OpenStudy (e.mccormick):

\[\frac{1-\tan ^{2}u}{1-\tan ^{2}u}=1\]And you can multiply anything by 1 and not change an equation.

OpenStudy (anonymous):

Ohhh! That makes sense. But there isn't a 1 in the equation.

OpenStudy (anonymous):

OOOOO...sorry. I'm not thinking.

OpenStudy (e.mccormick):

\[\tan 2x-\tan x=0\implies \frac{2\tan x}{1-\tan ^{2}}-\tan x=0\]What we need to change is that -tanx by multiplying it by 1.

OpenStudy (anonymous):

o ok

OpenStudy (e.mccormick):

So, if we have that special form of 1.... what would happen if we did a little multiplication?

OpenStudy (anonymous):

It would change the denominator to the common denominator.

OpenStudy (e.mccormick):

Exactly!

OpenStudy (anonymous):

Wow. I way overthink these problems.

OpenStudy (anonymous):

So, how do I proceed when I have everything common denominator-y.

OpenStudy (e.mccormick):

\[\frac{2\tan x}{1-\tan ^{2}x}-\tan x\cdot \frac{1-\tan ^{2}x}{1-\tan ^{2}x}=0\implies \frac{2\tan x}{1-\tan ^{2}x}- \frac{(\tan x)(1-\tan ^{2}x)}{1-\tan ^{2}x}=0\]

OpenStudy (e.mccormick):

Well, now that you can see it. What is next?

OpenStudy (anonymous):

I want to say...multiply the top of the right.

OpenStudy (anonymous):

Distribute the tangent of x.

OpenStudy (e.mccormick):

That looks good. Might have to factor it at some point though.... hmmm... I wonder if there is a factor by grouping hiding there somewhere.

OpenStudy (e.mccormick):

\[\frac{(2\tan x)-(\tan x)(1-\tan ^{2}x)}{1-\tan ^{2}x}\] \[\frac{\tan x[(2)-(1)(1-\tan ^{2}x)]}{1-\tan ^{2}x}\] \[\frac{\tan x[2-1+\tan ^{2}x]}{1-\tan ^{2}x}\] \[\frac{\tan x[1+\tan ^{2}x]}{1-\tan ^{2}x}\] Herm.

OpenStudy (e.mccormick):

Not quite what I expected... but I think it will work. Multiplying and combinding would have given the same thing.

OpenStudy (anonymous):

Ok.

OpenStudy (e.mccormick):

\[\frac{(2\tan x)-(\tan x)(1-\tan ^{2}x)}{1-\tan ^{2}x}\] \[\frac{2\tan x-\tan x+\tan ^{3}x}{1-\tan ^{2}x}\] \[\frac{\tan x+\tan ^{3}x}{1-\tan ^{2}x}\] And you can see how that ends up the same. either way is just as good.

OpenStudy (anonymous):

Do we have to do all this just to find the answers? Aren't we going to use the unit circle?

OpenStudy (e.mccormick):

Well, I was hoping to get rid of the bottom!

OpenStudy (anonymous):

Ok.

OpenStudy (e.mccormick):

So there is a little more work.... Cause I did not get the factors I expected. See, I was hoing the top would become something that canceled the bottom. =/

OpenStudy (anonymous):

I hate my teacher. Why would he give us 129 questions for our final and not be there for questions, and when you ask questions in the short time you DO see him, he says he can't help out?

OpenStudy (anonymous):

It's bad teaching.

OpenStudy (e.mccormick):

Fun. OK. So lets see... that bottom does factor. It is a difference of squares. Also, it can be replaced ... hmm... cotangent identitiy I think.

OpenStudy (e.mccormick):

Ah, secent... hmmm... might not be bad.

OpenStudy (e.mccormick):

Look at the Pythagoran identities.

OpenStudy (anonymous):

That won't work. The one you're thinking of is \[1+\tan ^{2}\theta=\sec ^{2}\]θ

OpenStudy (e.mccormick):

Yes.... but it might... I am trying something, which is an important thing in trig.

OpenStudy (anonymous):

Alrighty.

OpenStudy (e.mccormick):

I found something too ugly to contemplate. LOL! Want to see?